UT Sep 21 2022 - Night #12 of Restart




Night Info

  • Sunset 18:19
  • Nautical Twilight Ends 19:12
  • Astro Twilight Ends 19:41
  • Astro Twilight Begins 04:45
  • Nautical Twilight Begins 05:14
  • Sunrise 06:07
  • Moonrise 01:48
  • Moon illumination ~24%

Staff

  • OSA: D.G.Huerta
  • ISp: E. Solheid
  • Mtn Manager: R. Hannsen
  • ISA/SO: B.Rothberg, R.T.Gatto
  • Technical Expert: J.Hill
  • AO: D. Miller

Night Plan

ALTA prediction for Tonight

20220920_see_evol_time_Height_20_20000.png

MODS On-Sky Checkout

  • If MODS is released starting configuration is bino-MODS
  • If MODS unavailable (e.g. not cold engouh) but AGWs released - begin with AGW checkout, then move to Field Aberration and AGW1/2 Transform, then return to MODS
AGW 5/6 Checkout
  • Acquire on pointing star
  • Collimation check
  • Measure position of WFS hotspot in twilight
    • Check SH illumination
      • made backup copies of /home/telescope/TCS/Configuration/GCS/MODS_L.cfg andMODS_R.cfg
      • changed maxOffsetToHotspot and maxHotspotFromProbe to 2.54
      • Use the arrow controls in the GCSGUI eng panel
  • check MODS1/MODS2 guide probe wrt/rotator center
    • Files needed:
    • agw.ini located on mods1 and mods2 computers in: /home/mods/Config/MODS1 or /home/mods/Config/MODS2
MODS Verification
  • MODS Rotator Center measurement - high (0.5 hr)
    • Requires pointing mask (facility mask)
    • Take pairs of images 180 degrees apart
  • MODS AGW 5/6 - Measure Rotator zero point (mask alignment) - high (0.5 hr) - mask exchange required
    • Target Masks:
    • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1
      • The files needed are
      • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
  • MODS AGW Transform data -Meidum (High only if RotZP change) 1hrs
    • 1.2" seeing or better
    • If RotZP is > 0.1 deg then proceed with transform otherwise skip
      • if time, take the transform (after other tasks completed) - 12 points
    • must do RotZP First
    • Targets
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
  • Spectrophotometric standard all modes -High (to compare SX vs DX) - Critical - 1hr-2hr
    • If non-photometric - take 1 image and go through the modes in order of priority below, if photometric - 3 images
    • Pick any spectro-photometic standard star visible
      • Dual Grating
      • Red Grating
      • Blue Grating
      • Dual Prism
      • Red Prism
      • Blue Prism
  • Test ScienceOB - high 1hr
    • Any conditions
    • LS
      • Arp 230 (dual grating simple bright target - any conditions observable) - visible 22:30-04hr
        • poor weather science script already prepped w/finder chart
    • Imaging
      • Target?
      • Could be photometric fields if clear
    • If short on time - stop here and move to Field Aberration and AGW1/2 Transforms, then resume
  • Non-sidereal Guiding - Medium -1hr
    • Any condtions
    • imaging
      • Target?
    • LS
      • Target
Other MODS Checks
  • Lower priority relative to other MODS tests and taking AGW Transform confirimatory data and Field Aberration Confirmatory Data
  • Photometric Standards
    • any seeing, but must be clear
  • Skyflats - Medium
  • Async Offsets Test IT 8518 - 2hr Medium/High
    • binocular imaging program with radec offsets that are made asynchronously
      • small offsets
      • large offsets
    • binocular imaging program with detxy offsets that are made asynchronously
      • small offsets
      • large offsets
  • Mixed Mode LBC+MODS - 1hr Medium
    • poor weather acceptable
  • Mixed Mode MODS+LUCI - 1hr Medium
    • poor weather acceptable
  • Dither w/pointing mask
  • GCS and IDL have opposite field correction (IT 8315) - Medium/(1-1.5hrs)
    • Requires 1" or better
    • run IDL task field_collect (same as measuring field aberrations)
    • take LUCI data first to cross-check IDL vs GCS

Field Aberration Confirmatory Data

  • Reconfigure for bino-LUCIs
  • Conifrmatory Field Aberration (2 hrs) - Critical (1hour) - update installed already
  • Can be done in parallel (data collection, analysis in serial)
  • need both AGW1 and AGW2 on the telescope and checked to execute - DOES NOT NEED LUCIs - Just AGW 1 and AGW 2
  • 1.2" or better seeing (1" or better ideal)
  • Have OSA go to field closeby, point and collimate, then preset to target
  • Target Fields
    • BS9177 18:39:33.7 +49:05:38.0
    • StoneP1 22:37:45.0 -01:13:40.45
    • BS9188 23:30:33.4 +38:18:57.0
    • BS9107 03:32:03.0 +37:20:40.0

LUCI AGW1/2 Transform Confirmation

  • Collect transform data - (2hrs) - confirmatory - Requires LUCIs - transforms done and will be installed before sunset
    • update oaccontrol.conf with new results from D.Miller
      • update oaccontrol.conf with new results from D.Miller
      • backup file first
      • add the snippets from D.Miller in the correct place for AGW1 and AGW2
      • need to restart PCS and GCS to take hold
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
    • Can be done wth LUCI SW in crippled state (does not need flex comp, remember to load EmptyQueue.xml)

Backup Poor Weather

LUCI
  • LUCI Tasks - once poor weather MODS are done
    • LUCI Scripts
      • LUCI_2022_Restart_V1.81.xml - all scripts (except LS Rot ZP)
      • LUCI_Phot_ZP_OT_V6.xm - LUCI photometric standards (fixes AGW patrol problem)
    • Imaging check (1hr) - medium priority
    • ANY seeing OK
    • dither in-field
      • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am - only need to run 4 dithers
      • dither out-of-field (run one)
    • dither out-of-field (run one)
      • M52 - JHK, dither 5' out of field - visible all night
      • M40 - Broad filters, few narrow filters - visible all night

Night Log

Closed Dome
  • UT 00:28 - logged into OBS1 - brought up IIF, GCS guis, bringing up LUCI GUIs for later - holding off on any MODS stuff until O.Kuhn releases it.
  • UT 00:43 - taking test exposures
    • L1 0001-0002 - 2.51sec dark ok
    • L2 0001-0002 - 2.51sec dark ok
  • UT 01:54 - closed for humdiity and clouds - MODS checks
    • M2 BC offline - computer issues requiring J. Mason
    • MODS-1 mask issue - chute for loading mask into cassette seems bent, checking with OSU folks
    • both MODS still near room temperature - cooling is in progress - if we open - AGW5/6 checkout then switch to LUCIs
  • UT 02:54 - MODS are basically unavailable tonight - except for AGW 5/6 checkout - O.Kuhn will verify shortly that AGW cameras can be read out
    • in the morning E.Solheid will watch as masks are moved in the cassettes to make sure all ok
  • UT 03:13 - able to initialize and take images with AGW5 (guide and WFS)
  • UT 03:23 - still troubleshooting AGW6
  • UT 03:28 - restarted MODS2cam, then azcam server, then GCS, AGW6 can now readout images (guide and WFS)
  • UT 03:35 - need to reboot M1BC - caliban can't see the shared SCSI disks
  • UT 03:57 - can see some stars overhead, even the MW, but still thick clouds, and humidity 99.9% - Acquire preset successful closed dome (can check that off for MODS)
  • UT 04:56 - taking some LBC biases
  • UT 05:12 - O.Kuhn testing small update to DOFPIA w/askme
  • UT 06:06 - still closed for clouds and 99.9% humidity
  • UT 08:40 - still closed, can't see sky, humidity 99.9%
  • UT 10:05 - still closed, can't see sky, humidity 99.9%
  • UT 11:25 - same, see above
  • UT 12:42 - calling it

-- BarryRothberg - 19 Sep 2022
Topic revision: r7 - 21 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback